Charges of q+ and q- at the two vertices of an equilateral triangle. What can be said of the electric field E and potential V at the unoccupied vertex?

1
Expert's answer
2022-03-21T12:45:51-0400

Charges of q+ and q- at the two vertices of an equilateral triangle.

We can say of the electric field E and potential V


Due to charge q1 at point P

"E_1=\frac{kq_1}{r^2}=\frac{kq}{r^2}"

"E_2=\frac{kq_2}{r^2}=\frac{-kq}{r^2}"

"E_{net}=\sqrt{E_1^2+E_2^2+2E_1E_2cos\phi}"

"\phi=120°"

"E_{net}=\sqrt{E^2+E^2+E^2}=\sqrt{3}E"

"E_{net}=\sqrt{3}\times\frac{kq}{r^2}"

"E\not=0"

"V_{net}=V_1-V_2=0"

Result

"E\not=0,V=0"